OpenStudy (anonymous):

What is the means-to-MAD ratio of the two data sets, expressed as a decimal to the nearest tenth? Data Set 1: {85, 83, 90, 70, 87} Data Set 2: {75, 74, 68, 83, 80}

OpenStudy (anonymous):

we have to find the mean values of both data sets, first

OpenStudy (anonymous):

do you know the first 1?

OpenStudy (anonymous):

83

OpenStudy (anonymous):

yes good job

OpenStudy (anonymous):

do you know the second 1?

OpenStudy (anonymous):

75

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now, we have to compute the MAD of each data set

OpenStudy (anonymous):

for first data set, we have to compute these quantities: 85−83=...?83−83=...?90−83=...?70−83=...?87−83=...?

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

85−83=2 83−83=0 90−83=7 70−83= -13 87−83=4

OpenStudy (anonymous):

so the answer is 4?

OpenStudy (anonymous):

Now we have to consider the absolute values of those quantities, so the MAD for first data set is given by the subsequent computation: MAD = 2 + 0 + 7 + 13 + 4 / 5 + ...? MAD=2+0+7+13+45=...? please note that I have considered +13, and not -13, since we have to keep the absolute values

OpenStudy (anonymous):

26/5

OpenStudy (anonymous):

MAD=26/5=5.2

OpenStudy (anonymous):

then the ratio mean-to.-MAD, for first data set, is: 83/5.2=...?

OpenStudy (anonymous):

15.96153846

OpenStudy (anonymous):

15.96 rounded

OpenStudy (anonymous):

thank u